OpenStudy (anonymous):

If cos x=0.42 which of the following is cos (x+12pi) a) 0.042 b)0.04 c)0.54 d)0.42

OpenStudy (anonymous):

cosine is periodic with period \(2\pi\) so \[\cos(x)=\cos(x+2\pi)=\cos(x+4\pi)=\cos(x+6\pi)=...\]